OpenStudy (moongazer):

Determine the equivalent slope of the line based on the given conditions. 1. through points (-4, 8) and (2, -3) is it 11?

OpenStudy (anonymous):

\[m=\frac{y_2-y_1}{x_2-x_1}\]\[m=\frac{-3-8}{2+4}\]\[m=\frac{-11}{6}\]

OpenStudy (moongazer):

if not how do you find it also for whose equation is 3x – 4y = 12. whose equation is y = -3.

OpenStudy (moongazer):

Thanks! I forgot that the formula for slope is that. :)

OpenStudy (anonymous):

for 3x-4y=12 we need to write it in the form y=mx+c 3x-4y=12 -4y=12-3x y=-3+3/4x So the gradient m=3/4

OpenStudy (moongazer):

What is gradient?

OpenStudy (anonymous):

the slope

OpenStudy (moongazer):

other term for slope is gradient?

OpenStudy (anonymous):

correct

OpenStudy (moongazer):

Thanks for that information. :)

OpenStudy (anonymous):

For y=-3 use the formula of the slope again m=(y2-y1)/(x2-x1) =(-3+3)/(2-1) =0/1 =0

OpenStudy (anonymous):

we know that for y=-3 at any x value y=-3

OpenStudy (moongazer):

Can we just consider that m=0 in "For y=-3" that "m=0" because there is no x and that means that it is y=0x-3 sorry for my bad grammar.

OpenStudy (anonymous):

yes that's another way to do it, :D

OpenStudy (moongazer):

how about this? parallel to the line whose equation is x = 2.

OpenStudy (anonymous):

Because x=2 has a undetermined (m=some number/0) slope a parallel line would be x=3 or x=4 or x= any number

OpenStudy (moongazer):

Sorry but I don't get it. :)

OpenStudy (moongazer):

could explain it further?

OpenStudy (anonymous):

we can write x=2 as y*0=x-2 okay?

OpenStudy (moongazer):

I think I slightly got your explanation when I imagined a graph.

OpenStudy (moongazer):

anyway.. please continue to your explanation. :)

OpenStudy (anonymous):

and rewriting it in the form y=mx+c we get y=x/0-2/0 but x/0 is undefined so we need to find an equation that has an undefined slope. y=ax/0+c/0 where a and c are constants

OpenStudy (anonymous):

so y=ax/0+c/0 y=(ax+c)/0 y*0=(ax+c) ax+c=0 ax=-c x=-c/a

OpenStudy (anonymous):

does this make more sense?

OpenStudy (moongazer):

i'll just analyze it wait for a while. :)

OpenStudy (moongazer):

I didn't get it.

OpenStudy (moongazer):

could you graph it?? and btw what is the answer again. I might get it if I know the answer.

OpenStudy (anonymous):

 The answer is x=-c/a where c is any value and a can be any number but 0 So in the graph we have x=2, a line parallel to it is x=-3 and so is x=-1, x=9, x=11, ...., x=anynumber